# Floquet interferometry of a dressed semiconductor quantum dot

### v1

@@ II Coherent interaction of a dressed Quantum Dot with a reservoir
In order to probe this interaction, we can couple the dressed QD with a charge reservoir and dispersively measure its interaction with a radio-frequency (RF) cavity with resonant frequency $\omega_{\text{RF}}$ (Fig 1b). For a p-doped reservoir, such a configuration is commonly referred to in the literature as a single-hole box (SHB) [29]. In Ref. [30], we show how such a system can be effectively modelled as a driven two-level system (whose states are the occupied and unoccupied QD), obeying the time-dependent Hamiltonian
@@ II.2 Temperature Dependence
Figure 3: Effect of thermal broadening on Floquet interferometry: (a) energy-level schematic showing the reservoir and a Floquet ladder of QD + MW states. Thermal broadening produces a broadening of the reservoir Fermi surface. (b) Phase response of the DRT measured at three different mixing-chamber temperatures, as indicated. (c) FWHM extracted as a function of mixing-chamber temperature $T$. The three measurement points (10 mK, 210 mK and 620 mK) are indicated in blue, green and red, respectively. For reference, we also indicate the hole temperature $k_{\text{B}}T_{\text{h}}$ and the photon energy $\hbar\omega_{\text{MW}}$. (d-f) Experimental measurements of Floquet interferometry carried out with $\omega_{\text{MW}}/2\pi=27.5$ GHz at the three different temperatures. (g-h) Simulated response corresponding to the above measurements.
@@
$\text{FWHM}_{0}\left[V\right]=\frac{3.53k_{\text{B}}}{e\alpha}\sqrt{T_{MC}^{2}+T_{\text{h}}^{2}}.$ (7)
@@ II.3 Finite RF Probe Amplitude
Lastly, we explore the case where the RF probe is no longer small compared to the other broadening parameters ($k_{\text{B}}T_{\text{h}},h\Gamma$). This regime is particularly interesting from a theoretical perspective, as, if $\omega_{\text{RF}}$ and $\omega_{\text{MW}}$ are incommensurate, the Hamiltonian of the system is no longer periodic in time, so standard Floquet theory does not apply. Moreover, we can see how in Eq. (2) $H_{\text{RF}}$ is exactly identical to $H_{\text{MW}}$, and so we expect the RF and MW dressing of the QD to have a similar effect on the measurement. In Fig. 4, however, we see that this is clearly not the case. Here, we monitor the resonator response for increasing RF amplitude at two different $\delta\varepsilon_{\text{MW}}$ (Fig. 4c,e), observing a starkly different effect of the MW drive and the RF probe. The latter, in fact, rather than creating peaks at the different photon lines, separated by interference fringes, results in a single peak, which broadens linearly with $\delta\varepsilon_{\text{RF}}$, a phenomenon known as power broadening [30, 32, 29]. To understand the physical origin of this different response we can borrow a result from Appendix C and study the gate current produced by the (doubly dressed) QD exchanging particles with the reservoir, which can be written in the form
@@
|$\ket{\phi(t)}=$ $e^{i\varepsilon_{0}t/\hbar}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}J_{m}\left(\frac{\delta\varepsilon_{\text{MW}}}{\hbar\omega_{\text{MW}}}\right)$ (9)
@@
which resembles Eq. (1) but describes two interdigitated Floquet ladders, one for the RF and one for the MW. In particular, given the large difference in photon energy in the experiment, it is most useful to consider one RF ladder originating from each MW rung (Fig. 4a). We can now formally consider the interaction with the reservoir as Floquet scattering, causing transitions between different rungs [30, 36, 37, 38, 5, 39, 40, 41], where, as stated above, we are only sensitive to |$\ket{\phi_{m,n}}\rightarrow\ket{\phi_{m,n\pm 1}}$. Since we must only consider transitions with the same $m$ (i.e. $M=0$), we can still write the admittance of the QD as the sum in Eq. (5), only needing to substitute the small-signal response $Y_{0}$ with its power-broadened form
$Y_{PB}(\varepsilon_{0})=2\hbar\frac{(\alpha e)^{2}}{\delta\varepsilon_{\text{RF}}}\frac{\Gamma\omega_{\text{RF}}^{2}}{\omega_{\text{RF}}-i\Gamma}\sum_{n=1}^{+\infty}nJ_{n}^{2}\left(\frac{\delta\varepsilon}{\hbar\omega}\right)$ (10)
@@
where $\mathcal{F}(\varepsilon_{0})$ is the antiderivative of $\mathcal{F}^{\prime}(\varepsilon_{0})$, i.e. the convolution of the Fermi-Dirac of the Reservoir and the Lorentzian density of states of the QD. Notably, also Eq. (10) takes the form of a superposition of Bessel functions, similarly to the dressing caused by the MW, the difference with Eq. (5) originating from the fact that we are selecting transitions between adjacent RF rungs via the resonator.
Moreover, Eq. (10) highlights how, unlike the previous experiments, increasing the RF probe gives rise to a coherent source of broadening. This becomes apparent in Fig. 4e, where we apply $\delta\varepsilon_{\text{MW}}\approx 2\hbar\omega_{\text{MW}}$ to strongly suppress the $m=0$ rung in favour of the $m=\pm 1$ (purple line in Fig. 4b). Increasing the RF amplitude such that $\delta\varepsilon_{\text{RF}}>\hbar\omega_{\text{MW}}$ then causes the two rungs to interfere constructively, giving rise to an enhancement of the signal, in perfect agreement with our simulation.
Lastly, we note the absence of additional interference fringes in the superposition of the two power broadening fans, which have instead been observed elsewhere in the literature [42, 5, 43]. This is caused by the absence of any avoided crossing between charge states, which would lead to coherent accumulation of a diabatic phase via Landau-Zener-Stuckelberg-type processes. In our system, instead, the origin of the interference is caused by the different photon-number states in the Floquet rungs [5, 2], which can be pictured as a gap opening at the edge of the Floquet-Brillouin zone where different Floquet rungs become resonant [44, 2, 45]. Since $\omega_{\text{MW}}$ is incommensurable with $\omega_{\text{RF}}$, the separate (blue) $rf$ ladders originating from the (green) $MW$ rungs in Fig. 4a remain interdigitated, avoiding further Floquet resonances and thus the opening of additional gaps. This highlights the stochastic nature of the dispersive signal, as the reservoir incoherently exchanges holes QD. Yet, the quantum system shows signs of coherent interference between the two separate Floquet ladders.
@@ III Conclusion
Our work has investigated the non-equilibrium dynamics of a dressed hole QD stochastically exchanging charges with a reservoir. We have shown the emergence of interference fringes due to coherent multi-photon transitions, revealing the underlying structure of Floquet states. We reemphasize that the system lacks an avoided anticrossing, as is commonly observed in double QD systems, highlighting the novelty of our observation. Overall, this work demonstrates how, despite tunnelling in a reservoir completely randomizing the hole charge phase, this does not preclude quantum interference phenomena, allowing instead their dispersive sensing via the system’s electrical response. Having indicated the conditions under which single QD Floquet interferometry is observable, we predict the technique will be widely used for fast and accurate capacitance matrix extraction of these QD-reservoir systems. Particularly, the photon line spacing provides a direct measurement of the gate lever arm, that combined with cross capacitance measurements, reveals the full set of electrostatic parameters. The technique provides an alternative to experimentally time consuming methods like thermal broadening measurements [35] needed to fully characterise the system. Future measurements could be directed to exploring the role of multi-microwave excitations and the potential role of quantized density of states in the reservoir.
@@
Fabrication details. The transistor used in this study consists of a single gate silicon-on-insulator (SOI) nanowire transistor with a channel width of 120 nm, a length of 60 nm and height of 8 nm. The SOI consists of a 145-nm-thick buried oxide and a 8-nm-thick silicon layer with a Boron doping density of $5\cdot 10^{17}$  cm-3 . The silicon layer was patterned to create the channel using optical lithography, followed by a resist trimming process. The transistor gate stack consists of 1.9 nm HfSiON capped by 5 nm TiN and 50 nm polycrystalline silicon, leading to a total equivalent oxide thickness of 1.3 nm. After gate etching, a Si3N4 layer (10 nm) was deposited and etched to form a first spacer on the sidewalls of the gate, then 18-nm-thick Si raised source and drain contacts were selectively grown before source/drain extension implantation and activation annealing. A second spacer was formed, followed by source/drain implantations, an activation spike anneal and salicidation (NiPtSi). The nanowire quantum device and superconducting resonator were connected via aluminium bond wires.
[…]
@@ Appendix A Coherent interactions of a different quantum dot with a reservoir
Figure 5: (a) Stability diagram showing the Dot-Reservoir-Transition. Measurements are taken at the dashed line indicated. (b-d) Experimental data of Floquet interferometry carried out at three different frequencies as indicated. (e-h) Simulated response of the above interferometries.
To demonstrate the reproducibility of our observation, we repeat the experiment as described in Fig. 1 for a different QD in the same device, this time produced by a Boron dopant. In Fig. 5, we show the stability diagram of the transition showing a DRT intersected by several transitions with other QDs. We focus on the DRT indicated by the dashed line.
@@ Appendix B Coherent interactions of a quantum dot in a different device with a reservoir
Figure 6: (a) Stability diagram showing the Dot-Reservoir-Transition and nominal charge occupation N/N+1. Measurements are taken at the dashed line indicated. (b-g) Experimental data of Floquet interferometry carried out at six different frequencies as indicated.
To further illustrate the reproducible nature of our results, we repeat the experiment as described in Fig. 1 for a QD in a different device. In Fig. 6a, we show the stability diagram of the transition showing a DRT intersected by several transitions with other QDs. We focus on the DRT indicated by the dashed line.
@@
Appendix C Explicit Calculation of the Admittance of a Dressed SHB
In this section we will sketch the derivation of the high-frequency admittance of the doubly dressed Single Hole Box, effectively extending the formalism in Ref. [30] to the MW-driven case.
To describe the quantum dynamics of the SHB, we can picture it as composed of four subsystems: a single level in the QD, a charge reservoir (with temperature $T_{\text{h}}$), and two separate photon sources, the RF and the MW. Each component can be described by the Hamiltonians
$H_{QD}=\varepsilon_{0}~{}d^{\dagger}d$ (11)
$H_{R}=\sum_{\epsilon}~{}\epsilon~{}c_{\epsilon}^{\dagger}c_{\epsilon}$ (12)
@@ Appendix C Explicit Calculation of the Admittance of a Dressed SHB
where $d$, $c$, $a$ and $b$ represent the destruction operators of the QD, reservoir, RF, and MW, respectively. The interactions between them, furthermore, can be modelled as
$H_{QD-R}=\sum_{\epsilon}~{}Vc_{\epsilon}d^{\dagger}+V^{*}c_{\epsilon}^{\dagger}d$ (15)
$H_{QD-RF}=g_{\text{RF}}(a+a^{\dagger})d^{\dagger}d$ (16)
$H_{QD-MW}=g_{\text{MW}}(b+b^{\dagger})d^{\dagger}d,$ (17)
where we have considered the experimentally-relevant case of longitudinal QD-photon coupling. The parameter $g_{\text{RF}}$($g_{\text{MW}}$) is the coherent QD-RF (QD-MW) coupling, and $V_{\epsilon}$ is the QD-Reservoir coupling, which, in the wide-band approximation, we consider independent of the Reservoir energy $\epsilon$ [47, 48]. The Lindblad approximation is thus justified assuming that the thermalization times of a hole in the Reservoir is much faster than the total charge tunneling rate with the QD. The latter reads $\Gamma=|V|^{2}\mathcal{D}$, with $\mathcal{D}$ being the (flat) density of states in the Reservoir. The Hamiltonian naturally defines the (semiclassical) oscillation amplitude in terms of the time-evolution of the field operator, considered in a coherent state, according to [30, 31]
@@
and similarly for the MW. We shall note that in this work we always remain in the weak-coupling approximation between the QD and the radiation ($g\ll\hbar\omega$), so that the effect of spontaneous emission [30, 38, 49], as well as strong-coupling phenomena [43, 41, 50], can be neglected.
To establish links with Floquet theory, which we have relied upon in the main text, we can perform a Lang-Frisov transformation and study the problem in the polaron frame of reference [51, 52, 50]. This is achieved, defining the operator
@@
and considering the canonical transformation $e^{-S_{\text{RF}}}He^{S_{\text{RF}}}$, and similarly for the MW. It is important here to note that, as stressed in Section II.3, $[H_{QD-RF},H_{QD-MW}]=0$, and, thus $[S_{QD-RF},S_{QD-MW}]=0$. Therefore, the double dressing of the QD, or, more precisely, the two canonical transformations in the respective polaron frames can be considered interchangeably and without any additional interferences as $e^{S_{\text{RF}}}e^{S_{\text{MW}}}=e^{S_{\text{MW}}}e^{S_{\text{RF}}}=e^{S_{\text{RF}}+S_{\text{MW}}}$.
The Lindblad Master Equation can now be derived according to the standard theory [30, 53]. Notably, in Ref. [30] we show how the dynamics of the SHB can be considered without loss of generality as a two-level system, corresponding to the QD being empty or occupied. The unitary part of its dynamics is described by the Hamiltonian in Eq. (2), while the (stochastic) interaction with the reservoir can be described by the jump operators $L_{\pm}=(\sigma_{x}\pm i\sigma_{y})/2$ linked to the corresponding tunnel rates $\Gamma_{\pm}(t)$ tunneling in/out of the QD, with $\Gamma_{+}(t)+\Gamma_{-}(t)=\Gamma$ because of the conservation of charge. In this description, the probability of occupation of the QD satisfies
$\frac{d}{dt}P+\Gamma P=\Gamma_{-}(t),$ (20)
from which the gate current reads $I_{\text{g}}(t)=\alpha e\frac{d}{dt}P(t)$ and the admittance at the resonator frequency $\omega_{\text{RF}}/2\pi$ [30]
[…]
After tracing over the degrees of freedom not belonging to the QD, and the repeated application of Jacobi-Anger relation, we arrive to the results
$\Gamma_{-}(t)=\Gamma\sum_{m,M,n,N=-\infty}^{\infty}\tilde{J}_{n,N}\left(\frac{\delta\varepsilon_{\text{RF}}}{\hbar\omega_{\text{RF}}}\right)\tilde{J}_{m,M}\left(\frac{\delta\varepsilon_{\text{MW}}}{\hbar\omega_{\text{MW}}}\right)\mathcal{F}(\varepsilon_{0}+n\hbar\omega_{\text{RF}}+m\hbar\omega_{\text{MW}})\cos{(N\omega_{\text{RF}}t+M\omega_{\text{MW}}t)}$ (22)
@@
is the convolution of the Fermi-Dirac distribution of the reservoir and the Lorentzian effective density of states of the metastable QD (i.e. the antiderivative of Eq. (4)). Notably, already in the tunnel rate, we see the quantum mixing effect discussed in Section II.3. Moreover, we also see how, before the band-pass filtering of the resonator, the effect of the RF probe is indistinguishable from the MW dressing, as we would expect form the general theory. Combining Eq. (20) and Eq. (22), it is then trivial to find the gate current expressed in the form of Eq. (8), where we also have derived
@@
As a final remark, and for ease of replicating of the simulations presented in this work, we point out how it is computationally convenient to express the convolution between the Fermi-Dirac and the Lorentzian in terms of the polygamma functions, as
$\mathcal{F}^{\prime}(\varepsilon_{0})=\frac{1}{2}-\frac{1}{\pi}\Im\left[\psi_{1}\left(\frac{1}{2}+i\frac{\varepsilon_{0}}{2\pi k_{\text{B}}T_{\text{h}}}+\frac{h\Gamma}{2\pi k_{\text{B}}T}\right)\right],$ (27)
@@
$\mathcal{F}^{\prime}(\varepsilon_{0})=\frac{2}{\pi^{2}}\Re\left[\psi_{1}\left(\frac{1}{2}+i\frac{\varepsilon_{0}}{2\pi k_{\text{B}}T_{\text{h}}}+\frac{h\Gamma}{2\pi k_{\text{B}}T}\right)\right],$ (28)
@@
The Authors declare no Competing Financial or Non-Financial Interests
Data Avaliability

### v2

@@ II Coherent interaction of a dressed Quantum Dot with a reservoir
In order to probe this interaction, we can couple the dressed QD with a charge reservoir and dispersively measure its interaction with a radio-frequency (RF) cavity with resonant frequency $\omega_{\text{RF}}$ (Fig 1b). For a p-doped reservoir, this configuration is commonly referred to in the literature as a single-hole box (SHB) [29]. In Ref. [30], we show how such a system can be effectively modelled as a driven two-level system (whose states are the occupied and unoccupied QD), obeying the time-dependent Hamiltonian
@@ II.2 Temperature Dependence
Figure 3: Effect of thermal broadening on Floquet interferometry: (a) energy-level schematic showing the reservoir and a Floquet ladder of QD + MW states. Thermal broadening produces a broadening of the reservoir Fermi surface. (b) Phase response of the DRT measured at three different mixing-chamber temperatures, as indicated. (c) FWHM extracted as a function of mixing-chamber temperature $T_{MC}$. The three measurement points (10 mK, 210 mK and 620 mK) are indicated in blue, green and red, respectively. For reference, we also indicate the hole temperature $k_{\text{B}}T_{\text{h}}$ and the photon energy $\hbar\omega_{\text{MW}}$. (d-f) Experimental measurements of Floquet interferometry carried out with $\omega_{\text{MW}}/2\pi=27.5$ GHz at the three different temperatures. (g-h) Simulated response corresponding to the above measurements.
@@
$\text{FWHM}_{0}\left[V\right]=\frac{3.53k_{\text{B}}}{e\alpha}\sqrt{T_{\text{MC}}^{2}+T_{\text{h}}^{2}}.$ (7)
@@ II.3 Finite RF Probe Amplitude
Lastly, we explore the case where the RF probe is no longer small compared to the other broadening parameters ($k_{\text{B}}T_{\text{h}},h\Gamma$). This regime is particularly interesting from a theoretical perspective, as, if $\omega_{\text{RF}}$ and $\omega_{\text{MW}}$ are incommensurate, the Hamiltonian of the system is no longer periodic in time, so standard Floquet theory does not apply. Moreover, we can see how in Eq. (2) $H_{\text{RF}}$ is exactly identical to $H_{\text{MW}}$, and so we expect the RF and MW dressing of the QD to have similar effects on the measurement. In Fig. 4, however, we see that this is clearly not the case. Here, we monitor the resonator response for increasing RF amplitude at two different values of $\delta\varepsilon_{\text{MW}}$ (Fig. 4c,e), observing a starkly different effect of the MW drive and the RF probe. The latter, in fact, rather than creating peaks at the different photon lines, separated by interference fringes, results in a single peak, which broadens linearly with $\delta\varepsilon_{\text{RF}}$, a phenomenon known as power broadening [30, 32, 29]. To understand the physical origin of this different response we can borrow a result from Appendix C and study the gate current produced by the (doubly dressed) QD exchanging particles with the reservoir, which can be written in the form
@@
|$\ket{\phi(t)}=$ $e^{\text{i}\varepsilon_{0}t/\hbar}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}J_{m}\left(\frac{\delta\varepsilon_{\text{MW}}}{\hbar\omega_{\text{MW}}}\right)$ (9)
@@
which resembles Eq. (1) but describes two interdigitated Floquet ladders, one for the RF and one for the MW. In particular, given the large difference in photon energies in the experiment, it is most useful to consider one RF ladder originating from each MW rung (Fig. 4a). We can now formally consider the interaction with the reservoir as Floquet scattering, causing transitions between different rungs [30, 36, 37, 38, 5, 39, 40, 41], where, as stated above, we are only sensitive to |$\ket{\phi_{m,n}}\rightarrow\ket{\phi_{m,n\pm 1}}$. Since we must only consider transitions with the same $m$ (i.e., $M=0$), we can still write the admittance of the QD as the sum in Eq. (5), only needing to substitute the small-signal response $Y_{0}$ with its power-broadened form
$Y_{PB}(\varepsilon_{0})=2\hbar\frac{(\alpha e)^{2}}{\delta\varepsilon_{\text{RF}}}\frac{\Gamma\omega_{\text{RF}}^{2}}{\omega_{\text{RF}}-\text{i}\Gamma}\sum_{n=1}^{+\infty}nJ_{n}^{2}\left(\frac{\delta\varepsilon}{\hbar\omega}\right)$ (10)
@@
where $\mathcal{F}(\varepsilon_{0})$ is the antiderivative of $\mathcal{F}^{\prime}(\varepsilon_{0})$, i.e., the convolution of the Fermi-Dirac function of the reservoir and the Lorentzian density of states of the QD. Notably, also Eq. (10) takes the form of a superposition of Bessel functions, similarly to the dressing caused by the MW, the difference with Eq. (5) originating from the fact that we are selecting transitions between adjacent RF rungs via the resonator.
Moreover, Eq. (10) highlights how, unlike the previous experiments, increasing the RF probe gives rise to a coherent source of broadening. This becomes apparent in Fig. 4e, where we apply $\delta\varepsilon_{\text{MW}}\approx 2\hbar\omega_{\text{MW}}$ to strongly suppress the $m=0$ rung in favour of the $m=\pm 1$ (purple line in Fig. 4b). Increasing the RF amplitude such that $\delta\varepsilon_{\text{RF}}>\hbar\omega_{\text{MW}}$ then causes the two rungs to interfere constructively, giving rise to an enhancement of the signal, in complete agreement with our simulation.
Lastly, we note the absence of additional interference fringes in the superposition of the two power-broadening fans, which have instead been observed elsewhere in the literature [42, 5, 43]. This is caused by the absence of any avoided crossing between charge states, which would lead to coherent accumulation of a diabatic phase via Landau-Zener-Stückelberg-type processes. In our system, instead, the origin of the interference is caused by the different photon number states in the Floquet rungs [5, 2], which can be pictured as a gap opening at the edge of the Floquet-Brillouin zone where different Floquet rungs come on to resonance [44, 2, 45]. Since $\omega_{\text{MW}}$ is incommensurate with $\omega_{\text{RF}}$, the separate (blue) rf ladders originating from the (green) MW rungs in Fig. 4a remain interdigitated, avoiding further Floquet resonances and thus the opening of additional gaps. This highlights the stochastic nature of the dispersive signal, as the reservoir incoherently exchanges holes with the QD. Yet, the quantum system still shows signs of coherent interference between the two separate Floquet ladders.
@@ III Conclusion
Our work has investigated the non-equilibrium dynamics of a dressed hole QD stochastically exchanging charges with a reservoir. We have shown the emergence of interference fringes due to coherent multiphoton transitions, revealing the underlying structure of Floquet states. We re-emphasize that the system lacks an avoided anticrossing, as is commonly observed in double-QD systems, highlighting the novelty of our observation. Overall, this work demonstrates how, despite tunnelling to and from a reservoir completely randomizing the hole charge phase, this does not preclude quantum interference phenomena, allowing instead their dispersive sensing via the system’s electrical response. Having indicated the conditions under which single-QD Floquet interferometry is observable, we predict that the technique will be widely used for fast and accurate capacitance matrix extraction of these QD-reservoir systems. In particular, the photon line spacing provides a direct measurement of the gate lever arm, which, combined with cross-capacitance measurements, reveals the full set of electrostatic parameters. The technique provides an alternative to the experimentally time-consuming methods such as thermal-broadening measurements [35] needed to fully characterise the system. Future measurements could be directed to exploring the role of multi-microwave excitations and the potential role of a quantized density of states in the reservoir.
@@
Fabrication details. The transistor used in this study consists of a single-gate silicon-on-insulator (SOI) nanowire transistor with a channel width of 120 nm, a length of 60 nm and height of 8 nm. The SOI consists of a buried oxide 145 nm thick and an 8 nm silicon layer with a boron doping density of $5\cdot 10^{17}$ cm-3. The silicon layer was patterned to create the channel using optical lithography, followed by a resist-trimming process. The transistor gate stack consists of 1.9 nm HfSiON capped by 5 nm TiN and 50 nm polycrystalline silicon, leading to a total equivalent oxide thickness of 1.3 nm. After gate etching, a Si3N4 layer (10 nm) was deposited and etched to form a first spacer on the sidewalls of the gate, then 18-nm Si raised source and drain contacts were selectively grown before source/drain extension implantation and activation annealing. A second spacer was formed, followed by source/drain implantations, an activation spike anneal and salicidation (NiPtSi). The nanowire quantum device and superconducting resonator were connected via aluminium bond wires.
Measurement set-up. Measurements were performed at the base temperature of a dilution refrigerator ($T\sim 10$ mK). Low-frequency signals ($V_{\text{g}}$, $V_{\text{bg}}$) were applied through cryogenic filters, while radio-frequency and high-frequency microwave tones were applied through filtered coaxial lines to a coupling capacitor connected to the RF resonator, or through an on-PCB (printed circuit-board) bias-T connected to the Source of the transistor, respectively. The MW tones were produced by an Anritsu MG3694C Signal Generator. The resonator consists of a NbTiN superconducting spiral inductor ($L\sim 30$ nH), coupling capacitor ($C_{c}\sim 40$ fF) and low-pass filter fabricated by Star Cryoelectronics. For exact details see Ref. [46]. The PCB was made from RO4003C 0.8 mm thick with an immersion silver finish. The reflected RF signal was amplified at 4 K and room temperature, followed by quadrature demodulation (Polyphase Microwave AD0540B), from which the amplitude and phase of the reflected signal were obtained (homodyne detection).
@@ Appendix A Coherent interactions of a different quantum dot with a reservoir
Figure 5: (a) Stability diagram showing the dot–reservoir transition. Measurements are taken at the dashed line indicated. (b-d) Experimental data of Floquet interferometry carried out at the three different frequencies indicated. (e-h) Simulated response matching the above experimental results.
To demonstrate the reproducibility of our observations, we repeat the experiment as described for a different QD in the same device, this time produced by a boron dopant (see Fig. 1). In Fig. 5, we show the stability diagram of the transition showing a DRT intersected by several transitions with other QDs. We focus on the DRT indicated by the dashed line.
@@ Appendix B Coherent interactions of a quantum dot in a different device with a reservoir
Figure 6: (a) Stability diagram showing the dot–reservoir transition and nominal charge occupation $N$ or $N+1$. Measurements are taken at the dashed line indicated. (b-g) Experimental data of Floquet interferometry carried out at the six different frequencies indicated.
To further illustrate the reproducible nature of our results, we repeat the experiment for a QD in a different device (see Fig. 1). In Fig. 6a, we show the stability diagram of the transition, showing a DRT intersected by several transitions with other QDs. We focus on the DRT indicated by the dashed line.
@@
Appendix C Explicit Calculation of the Admittance of a Dressed Single-hole Box
In this section we sketch the derivation of the high-frequency admittance of the doubly dressed single-hole box (SHB), effectively extending the formalism in Ref. [30] to the MW-driven case.
To describe the quantum dynamics of the SHB, we can picture it as composed of four subsystems: a single level in the QD, a charge reservoir (with temperature $T_{\text{h}}$), and two separate photon sources, the RF and the MW. Each component can be described by a Hamiltonian,
$H_{\text{QD}}=\varepsilon_{0}~{}d^{\dagger}d$ (11)
$H_{\text{R}}=\sum_{\epsilon}~{}\epsilon~{}c_{\epsilon}^{\dagger}c_{\epsilon}$ (12)
@@ Appendix C Explicit Calculation of the Admittance of a Dressed Single-hole Box
where $d$, $c$, $a$ and $b$ represent the destruction operators of the QD, reservoir, RF and MW, respectively. The interactions between them, furthermore, can be modelled as
$H_{\text{QD-R}}=\sum_{\epsilon}~{}Vc_{\epsilon}d^{\dagger}+V^{*}c_{\epsilon}^{\dagger}d$ (15)
$H_{\text{QD-RF}}=g_{\text{RF}}(a+a^{\dagger})d^{\dagger}d$ (16)
$H_{\text{QD-MW}}=g_{\text{MW}}(b+b^{\dagger})d^{\dagger}d,$ (17)
where we have considered the experimentally relevant case of longitudinal QD-photon coupling. The parameter $g_{\text{RF}}$($g_{\text{MW}}$) is the coherent QD-RF (QD-MW) coupling, and $V_{\epsilon}$ is the QD-reservoir coupling, which, in the wide-band approximation, we consider independent of the reservoir energy $\epsilon$ [47, 48]. The Lindblad approximation is thus justified, assuming that the thermalization time of a hole in the reservoir is much faster than the total charge tunneling rate to or from the QD. The latter reads $\Gamma=|V|^{2}\mathcal{D}$, with $\mathcal{D}$ being the (flat) density of states in the reservoir. The Hamiltonian naturally defines the (semiclassical) oscillation amplitude in terms of the time evolution of the field operator, considered in a coherent state, according to [30, 31],
@@
and similarly for the MW. We shall note that in this work we always remain in the regime of weak coupling between the QD and the radiation ($g\ll\hbar\omega$), so that the effect of spontaneous emission [30, 38, 49], as well as strong-coupling phenomena [43, 41, 50], can be neglected.
To establish links with Floquet theory, which we have relied upon in the main text, we can perform a Lang-Frisov transformation and study the problem in the polaron frame of reference [51, 52, 50]. This is achieved by defining the operator
@@
and considering the canonical transformation $e^{-S_{\text{RF}}}He^{S_{\text{RF}}}$, and similarly for the MW. It is important here to note that, as stressed in Section II.3, $[H_{\text{QD-RF}},H_{\text{QD-MW}}]=0$, and, thus $[S_{\text{QD-RF}},S_{\text{QD-MW}}]=0$. Therefore, the double dressing of the QD, or, more precisely, the two canonical transformations in the respective polaron frames, can be considered interchangeably and without any additional interferences as $e^{S_{\text{RF}}}e^{S_{\text{MW}}}=e^{S_{\text{MW}}}e^{S_{\text{RF}}}=e^{S_{\text{RF}}+S_{\text{MW}}}$.
The Lindblad master equation can now be derived according to the standard theory [30, 53]. Notably, in Ref. [30] we show how the dynamics of the SHB can be considered without loss of generality as a two-level system, corresponding to the QD being empty or occupied. The unitary part of its dynamics is described by the Hamiltonian in Eq. (2), while the (stochastic) interaction with the reservoir can be described by the jump operators $L_{\pm}=(\sigma_{x}\pm\text{i}\sigma_{y})/2$ linked to the corresponding tunnel rates $\Gamma_{\pm}(t)$ tunneling in/out of the QD, with $\Gamma_{+}(t)+\Gamma_{-}(t)=\Gamma$ because of the conservation of charge. In this description, the probability of occupation of the QD satisfies
$\frac{\text{d}}{\text{d}t}P+\Gamma P=\Gamma_{-}(t),$ (20)
from which the gate current reads $I_{\text{g}}(t)=\alpha e\frac{\text{d}}{\text{d}t}P(t)$ and the admittance at the resonator frequency $\omega_{\text{RF}}/2\pi$ [30]
$Y=\frac{\alpha e\omega_{\text{RF}}}{\pi\delta\varepsilon_{\text{RF}}}\int_{0}^{\frac{2\pi}{\omega_{\text{RF}}}}e^{\text{i}\omega_{\text{RF}}t^{\prime}}I_{\text{g}}(t^{\prime})\text{d}t^{\prime}.$ (21)
After tracing over the degrees of freedom not belonging to the QD, and the repeated application of Jacobi-Anger relation, we arrive at the results
$\Gamma_{-}(t)=\Gamma\sum_{m,M,n,N=-\infty}^{\infty}\tilde{J}_{n,N}\left(\frac{\delta\varepsilon_{\text{RF}}}{\hbar\omega_{\text{RF}}}\right)\tilde{J}_{m,M}\left(\frac{\delta\varepsilon_{\text{MW}}}{\hbar\omega_{\text{MW}}}\right)\mathcal{F}(\varepsilon_{0}+n\hbar\omega_{\text{RF}}+m\hbar\omega_{\text{MW}})\cos{(N\omega_{\text{RF}}t+M\omega_{\text{MW}}t)},$ (22)
@@
is the convolution of the Fermi-Dirac distribution of the reservoir and the Lorentzian effective density of states of the metastable QD (i.e., the antiderivative of Eq. (4)). Notably, already in the tunnel rate, we see the quantum mixing effect discussed in Section II.3. Moreover, we also see how, before the band-pass filtering of the resonator, the effect of the RF probe is indistinguishable from the MW dressing, as we would expect form the general theory. Combining Eq. (20) and Eq. (22), it is then trivial to find the gate current expressed in the form of Eq. (8), where we also have derived
@@
As a final remark, and for ease of replicating of the simulations presented in this work, we point out how it is computationally convenient to express the convolution between the Fermi-Dirac function and the Lorentzian in terms of the polygamma functions, as
$\mathcal{F}(\varepsilon_{0})=\frac{1}{2}-\frac{1}{\pi}\Im\left[\psi_{0}\left(\frac{1}{2}+\text{i}\frac{\varepsilon_{0}}{2\pi k_{\text{B}}T_{\text{h}}}+\frac{h\Gamma}{2\pi k_{\text{B}}T}\right)\right],$ (27)
@@
$\mathcal{F}^{\prime}(\varepsilon_{0})=\frac{2}{\pi^{2}}\Re\left[\psi_{1}\left(\frac{1}{2}+\text{i}\frac{\varepsilon_{0}}{2\pi k_{\text{B}}T_{\text{h}}}+\frac{h\Gamma}{2\pi k_{\text{B}}T}\right)\right],$ (28)
@@
The authors declare no competing financial or non-financial interests.
Data Availability